An oscillatory motion has a period of 50.0s and an amplitude of 20.0cm. Determine its maximum acceleration and mechanical energy
1
Expert's answer
2021-01-31T03:39:04-0500

a)

"a_{max}=-A\omega^2=-A\dfrac{4\pi^2}{T^2},""a_{max}=-0.2\ m\cdot\dfrac{4\pi^2}{(50.0\ s)^2}=-3.16\cdot10^{-3}\ \dfrac{m}{s^2}."


The magnitude of the maximum acceleration is "3.16\cdot10^{-3}\ \dfrac{m}{s^2}".

b) By the definition of the law of conservation of energy, we have:


"ME=U+K=\dfrac{1}{2}kx^2+\dfrac{1}{2}mv^2."

An oscillatory motion of the oscillator is defined by the position:


"x(t)=Acos(\omega t+\phi)"

and velocity:


"v(t)=-A\omega cos(\omega t+\phi)."

Then, we can write:


"ME=\dfrac{1}{2}kA^2cos^2(\omega t+\phi)+\dfrac{1}{2}mA^2sin^2(\omega t+\phi)."

Applying trigonometric identity "cos^2\theta+sin^2\theta=1" and "\omega=\sqrt{\dfrac{k}{m}}", we get:


"ME=\dfrac{1}{2}kA^2=\dfrac{1}{2}\cdot k\cdot(0.2\ m)^2=0.02\ m^2\cdot k."